# Generalized finite and affine $W$-algebras in type $A$

### v1

@@ 4.2 The algebras $U(\lambda,\mu)$ and $Zhu_{H}W^{k}(\lambda,\mu)$
where $\phi_{\mathfrak{n}_{\lambda,\mu}}$, $\phi^{\mathfrak{n}_{\lambda,\mu}^{\ast}}$ are the odd vector superspaces defined in (3.3). The Poisson bracket on $\mathcal{C}(\lambda,\mu)$ is induced from the Lie bracket on $\mathfrak{a}$ and the super-commutator $[\phi^{\mathsf{m}},\phi_{\mathsf{n}}]=\mathsf{m}(\mathsf{n}).$ Consider the element
@@
$H(\mathcal{C}(\lambda,\mu),\mathcal{Q})\cong H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}}),$ (4.5)
@@
$\text{gr}(J^{cl}_{a})=\left(j_{a}-\frac{1}{2},-j_{a}+\frac{1}{2}\right),\quad\text{gr}(\phi^{I})=\left(\text{deg}(E_{I})+\frac{1}{2},-\text{deg}(E_{I})+\frac{1}{2}\right)$
@@
    $H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})=H^{0}(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})$.
@@
    Similarly to Lemma 3.9 (3), we can show that $H^{pq}(\text{gr}(\widetilde{\mathcal{C}}(\lambda,\mu)),\widetilde{\mathcal{Q}}^{gr})=0$ unless $p+q=0$, which means $\mathcal{Q}$ is a good differential. Moreover, since $\mathcal{Q}$ preserves the conformal weight and the subspace of $\widetilde{C}(\lambda,\mu)$ with a given conformal weight is finite dimensional, we can apply [6, Lemma 4.2].
@@
    As in Lemma 3.9 (2), the differential $\widetilde{\mathcal{Q}}$ is almost linear. Moreover, recalling notation (4.8), observe that since $H^{pq}(r_{-},\widetilde{\mathcal{Q}}^{\text{gr}})$ is nonzero only when $p+q=0,$ the cohomology $H(\text{gr}(\widetilde{\mathcal{C}}(\lambda,\mu)),\widetilde{\mathcal{Q}}^{\text{gr}})\simeq S(H(r_{-},\widetilde{\mathcal{Q}}^{\text{gr}}))$ is also concentrated at degree 00 part. Now, by part (1) we get part (2).
@@
The cohomology $H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})$ is a Poisson subalgebra of $S(\mathfrak{p}).$
@@
By the same argument as in the proof of Lemma 3.4, we can show $H(\mathcal{C}(\lambda,\mu),\mathcal{Q})$ has the Poisson algebra structure induced from that of $\mathcal{C}(\lambda,\mu)$. Since $H(\mathcal{C}(\lambda,\mu),\mathcal{Q})\simeq H^{0}(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}}),$ this is a Poisson subalgebra of $S(J_{\mathfrak{p}}^{cl}).$ Finally, since $J_{\mathfrak{p}}^{cl}\simeq\mathfrak{p}$ as Lie algebras, the proof is complete. ∎
Now we want to show that the Poisson algebra $H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})$ can also be realized in two different ways:
@@
Consider the Kazhdan filtration $K_{0}(\mathfrak{a})\subset K_{1}(\mathfrak{a})\subset K_{2}(\mathfrak{a})\subset\dots$ on $U(\mathfrak{a})$, where a homogeneous element $a\in\mathfrak{a}$ is in $K_{s}(\mathfrak{a})$ if and only if $a\in\mathfrak{a}(j_{a})$ for some $j_{a}\geqslant 1-s.$ Then for elements $a\in K_{s}(\mathfrak{a})$ and $b\in K_{t}(\mathfrak{a}),$ the commutator $ab-ba$ belongs to $K_{s+t-1}(\mathfrak{a}).$ Then the graded algebra $S(\mathfrak{a}):=\text{gr}^{K}(U(\mathfrak{a}))$ is the Poisson algebra endowed with the bracket
@@
Then $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$ and $S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu}$ are $\mathfrak{n}_{\lambda,\mu}$-modules via adjoint actions. Here we note that the Kazhdan filtration is the analogue of the filtration (4.7) induced from the conformal weight. In the following proposition, we describe the two realizations of $H(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{Q})$ in (4.9) more precisely via the $\mathfrak{n}_{\lambda,\mu}$-module $S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu}$.
@@
We show that the Chevalley–Eilenberg complex of the $\mathfrak{n}_{\lambda,\mu}$-module $S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu}$ is isomorphic to the complex $(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})$. Consequently, the cohomology of $(\widetilde{\mathcal{C}}(\lambda,\mu),\widetilde{\mathcal{Q}})$ is concentrated in degree 00, so the cohomology $H(\mathfrak{n}_{\lambda,\mu},S(\mathfrak{a})/\mathcal{I}^{cl}_{\lambda,\mu})$ is also concentrated in degree 00. By (4.6), the map between the complexes
@@
The Lie algebra cohomology of $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$ is concentrated on degree 00 part; that is,
@@
The main idea is the same as for the proof of [9, Proposition 5.2]. Consider the cochain complex
@@
where $\mathbb{C}_{-\chi}$ is the one-dimensional representation of $\mathfrak{n}_{\lambda,\mu}$ with $\mathsf{n}\cdot 1=-\chi(\mathsf{n})$. Hence we can follow the argument of [6, Appendix]. ∎
@@ 5.1 Principal nilpotent case
with $10\leqslant r\leqslant\lambda_{n-m+1}-$ 1 is a basis of $(\text{ker}\,\varphi)[m].$ Combining Corollary 3.11 and (5.1), we can derive following properties.
@@
$[w_{3}{}_{\lambda}w_{3}]=(\partial+2\lambda)\left(-(k+4)w_{4}+\left(\frac{k}{4}+1\right):w_{2}w_{2}:\right)$
and all other $\lambda$-brackets are 00. Note that, in this example, we can see that $W^{k}(\lambda,(n))$ does not have a conformal vector; see Remark 3.7. If we look for a conformal vector of weight $2$ in the form
@@
$[w{}_{\lambda}w]=(\partial+2\lambda)w+c\lambda^{3}$
@@
Then the following is an algebraically independent generating set of the center of the algebra $U(\mathfrak{a})$ [13, Cor. 2.7]:
@@
It is obvious that $\Psi_{m}^{(r)}\in U(\lambda,(n))$ since $[\mathsf{n},\Phi_{m}^{(r)}]=0$ for any $\mathsf{n}\in\mathfrak{n}_{\lambda,(n)}.$ Also, it is not hard to check that $\Psi_{m}^{(r)}$ is nonzero element in the quotient
@@
¯$\overline{\mathcal{M}}_{3}=\left[\begin{array}[]{ccc}x+2\lambda_{1}+\epsilon_{11}(u)&-u^{\lambda_{2}-1}&0\\ \epsilon_{21}(u)&x+\lambda_{2}+\epsilon_{22}(u)&-u^{\lambda_{3}-1}\\ \epsilon_{31}(u)&\epsilon_{32}(u)&x+\epsilon_{33}(u)\end{array}\right].$ (5.6)
We substituted $\epsilon_{12}(u)$, $\epsilon_{23}(u)$ and $\epsilon_{13}(u)$ by their respective images $-u^{\lambda_{2}-1}$, $-u^{\lambda_{3}-1}$ and 00 in $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}.$ Since the graded algebra of $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$ is commutative, we can find $\text{gr}(\Psi_{m}^{(r)})$ by computing the column determinant of (5.6) which is given by
@@ 5.2 Minimal nilpotent case
In this subsection, we describe generators of generalized finite and affine $W$-algebras when $\lambda$ is arbitrary and $\mu=(1,1,\dots,1,2)$ (with $n-2$ parts equal to $1$). From the $\mathbb{Z}$-grading (2.6),
@@
To apply Corollaries 3.11 and 4.7, we first describe the kernel of $\varphi$. Since $\chi=(E_{n-1,n}^{(\lambda_{n}-1)})^{\ast}$, we have
@@
$\mathcal{B}_{1}$ :=$:=\{E_{ij}^{(r)}\ |\ 1\leqslant i\leqslant n-2,1\leqslant j\leqslant n-1,\lambda_{j}-\text{min}(\lambda_{i},\lambda_{j})\leqslant r<\lambda_{j}\}$
[…]
@@
    $G_{n\alpha}^{(r)}:=J_{E_{n\alpha}^{(r)}}$ $+\sum_{a+b=\lambda_{n}-1+r}:J_{E_{n-1,\alpha}^{(a)}}J_{E_{nn}^{(b)}}:-\sum_{\gamma=1}^{n-2}\ \sum_{a+b=\lambda_{n}-1+r}:J_{E_{n-1,\gamma}^{(a)}}J_{E_{\gamma\alpha}^{(b)}}:$ (5.8)
@@
    By Theorem 3.10, it is sufficient to show that the above elements are inside $\widetilde{Q}(J_{a})=0$ for $a\in\mathcal{B}_{1}$. Let $21\leqslant i\leqslant n-$ 2 and $11\leqslant j\leqslant n-$ 1. Then
@@
$\mathcal{B}_{1}$ =$=\{E_{11}^{(0)},E_{12}^{(0)},E_{13}^{(1)},E_{21}^{(0)},E_{22}^{(0)},E_{23}^{(1)}\}\cup\{E_{33}^{(0)}+E_{44}^{(0)},E_{33}^{(1)}+E_{44}^{(1)}\}\ ,$
$\mathcal{B}_{2}$ =$=\{E_{41}^{(0)},E_{42}^{(0)},E_{43}^{(1)},E_{43}^{(0)}\}\ .$
For $a\in\mathcal{B}_{1}$, $a$ and $J_{a}$ are generators of $U(\lambda,\mu)$ and $W^{k}(\lambda,\mu)$, respectively, of conformal weight $1$. The generators of $U(\lambda,\mu)$ of conformal weight $2$ are
@@
If $\lambda=(1,1,2,3)$ and $\mu=(1,1,2)$, then the last terms of (5.7) and (5.8) are equal to 00. In this case, the generators of $W^{k}(\lambda,\mu)$ of conformal weight $2$ are

### v2

@@ 4.2 The algebras $U(\lambda,\mu)$ and $Zhu_{H}W^{k}(\lambda,\mu)$
where $\phi_{\mathfrak{n}_{\lambda,\mu}}$, $\phi^{\mathfrak{n}_{\lambda,\mu}^{\ast}}$ are the odd vector superspaces defined in (3.3). The Poisson bracket on $\mathcal{C}(\lambda,\mu)$ is induced from the Lie bracket on $\mathfrak{a}$ and the super-commutator $\phi^{\mathsf{m}},\phi_{\mathsf{n}}]=\mathsf{m}(\mathsf{n}).$ Consider the element
@@
[…]
@@
$\text{gr}(J^{cl}_{a})=\left(j_{a}-\frac{1}{2},-j_{a}+\frac{1}{2}\right),\quad\text{gr}(\phi^{I})=\left(-\text{deg}(E_{I})+\frac{1}{2},\text{deg}(E_{I})+\frac{1}{2}\right)$
@@
[…]
@@
    Similarly to Lemma 3.10 (3), we can show that $H^{pq}(\text{gr}(\widetilde{\mathcal{C}}(\lambda,\mu)),\widetilde{\mathcal{Q}}^{gr})=0$ unless $p+q=0$, which means $\mathcal{Q}$ is a good differential. Moreover, since $\mathcal{Q}$ preserves the conformal weight and the subspace of $\widetilde{C}(\lambda,\mu)$ with a given conformal weight is finite dimensional, we can apply [9, Lemma 4.2].
@@
    As in Lemma 3.10 (2), the differential $\widetilde{\mathcal{Q}}$ is almost linear. Moreover, recalling notation (4.8), observe that since $H^{pq}(r_{-},\widetilde{\mathcal{Q}}^{\text{gr}})$ is nonzero only when $p+q=0,$ the cohomology $H(\text{gr}(\widetilde{\mathcal{C}}(\lambda,\mu)),\widetilde{\mathcal{Q}}^{\text{gr}})\simeq S(H(r_{-},\widetilde{\mathcal{Q}}^{\text{gr}}))$ is also concentrated at degree 0 part. Now, by part (1) we get part (2).
@@
[…]
@@
[…]
@@
Consider the Kazhdan filtration $K_{0}(\mathfrak{a})\subset K_{1}(\mathfrak{a})\subset K_{2}(\mathfrak{a})\subset\dots$ on $U(\mathfrak{a})$, where a homogeneous element $a\in\mathfrak{a}$ is in $K_{s}(\mathfrak{a})$ if and only if $a\in\mathfrak{a}(j_{a})$ for some $j_{a}\geqslant 1-s.$ Then for elements $a\in K_{s}(\mathfrak{a})$ and $b\in K_{t}(\mathfrak{a}),$ the commutator abbaab-ba belongs to $K_{s+t-1}(\mathfrak{a}).$ Then the graded algebra $S(\mathfrak{a}):=\text{gr}^{K}(U(\mathfrak{a}))$ is the Poisson algebra endowed with the bracket
@@
[…]
@@
[…]
@@
The Lie algebra cohomology of $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$ is concentrated on degree 0 part; that is,
@@
The main idea is the same as for the proof of [17, Proposition 5.2]. Consider the cochain complex
@@
where $\mathbb{C}_{-\chi}$ is the one-dimensional representation of $\mathfrak{n}_{\lambda,\mu}$ with $\mathsf{n}\cdot 1=-\chi(\mathsf{n})$. Hence we can follow the argument of [9, Appendix]. ∎
@@ 5.1 Principal nilpotent case
with $0\leqslant r\leqslant\lambda_{n-m+1}-1$ is a basis of $(\text{ker}\,\varphi)[m].$ Combining Corollary 3.13 and (5.1), we can derive following properties.
@@
$w_{3}{}_{\lambda}w_{3}]=(\partial+2\lambda)\left(-(k+4)w_{4}+\left(\frac{k}{4}+1\right):w_{2}w_{2}:\right)$
and all other $\lambda$-brackets are 0. Note that, in this example, we can see that $W^{k}(\lambda,(n))$ does not have a conformal vector; see Remark 3.8. If we look for a conformal vector of weight $2$ in the form
@@
$w{}_{\lambda}w]=(\partial+2\lambda)w+c\lambda^{3}$
@@
Then the following is an algebraically independent generating set of the center of the algebra $U(\mathfrak{a})$ [23, Cor. 2.7]:
@@
It is obvious that $\Psi_{m}^{(r)}\in U(\lambda,(n))$ since $\mathsf{n},\Phi_{m}^{(r)}]=0$ for any $\mathsf{n}\in\mathfrak{n}_{\lambda,(n)}.$ Also, it is not hard to check that $\Psi_{m}^{(r)}$ is nonzero element in the quotient
@@
$\overline{\mathcal{M}}_{3}=\left[\begin{array}[]{ccc}x+2\lambda_{1}+\epsilon_{11}(u)&-u^{\lambda_{2}-1}&0\\ \epsilon_{21}(u)&x+\lambda_{2}+\epsilon_{22}(u)&-u^{\lambda_{3}-1}\\ \epsilon_{31}(u)&\epsilon_{32}(u)&x+\epsilon_{33}(u)\end{array}\right].$ (5.6)
We substituted $\epsilon_{12}(u)$, $\epsilon_{23}(u)$ and $\epsilon_{13}(u)$ by their respective images $u^{\lambda_{2}-1}$, $u^{\lambda_{3}-1}$ and 0 in $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}.$ Since the graded algebra of $U(\mathfrak{a})/\mathcal{I}_{\lambda,\mu}$ is commutative, we can find $\text{gr}(\Psi_{m}^{(r)})$ by computing the column determinant of (5.6) which is given by
@@ 5.2 Minimal nilpotent case
In this subsection, we describe generators of generalized finite and affine $W$-algebras when $\lambda$ is arbitrary and $\mu=(1,1,\dots,1,2)$ (with n2n-2 parts equal to $1$). From the $\mathbb{Z}$-grading (2.6),
@@
To apply Corollaries 3.13 and 4.7, we first describe the kernel of $\varphi$. Since $\chi=(E_{n-1,n}^{(\lambda_{n}-1)})^{\ast}$, we have
@@
$\mathcal{B}_{1}$ $=\{E_{ij}^{(r)}\ |1\leqslant i\leqslant n-2,1\leqslant j\leqslant n-1,\lambda_{j}-\text{min}(\lambda_{i},\lambda_{j})\leqslant r<\lambda_{j}\}$
$\hskip 170.71652pt\cup\{E_{n-1,n-1}^{(r)}+E_{nn}^{(r)}\ |0\leqslant r<\lambda_{n}-1\},$
$\mathcal{B}_{2}$ $=\{E_{n\alpha}^{(r)}\ |1\leqslant\alpha\leqslant n-1,0\leqslant r<\lambda_{\alpha}-1\}.$
@@
    $G_{n\alpha}^{(r)}=J_{E_{n\alpha}^{(r)}}$ $+\sum_{a+b=\lambda_{n}-1+r}:J_{E_{n-1,\alpha}^{(a)}}J_{E_{nn}^{(b)}}:-\sum_{\gamma=1}^{n-2}\ \sum_{a+b=\lambda_{n}-1+r}:J_{E_{n-1,\gamma}^{(a)}}J_{E_{\gamma\alpha}^{(b)}}:$ (5.8)
@@
    By Theorem 3.12, it is sufficient to show that the above elements are inside $\widetilde{Q}(J_{a})=0$ for $a\in\mathcal{B}_{1}$. Let $1\leqslant i\leqslant n-2$ and $1\leqslant j\leqslant n-1$. Then
@@
$\mathcal{B}_{1}$ $=\{E_{11}^{(0)},E_{12}^{(0)},E_{13}^{(1)},E_{21}^{(0)},E_{22}^{(0)},E_{23}^{(1)}\}\cup\{E_{33}^{(0)}+E_{44}^{(0)},E_{33}^{(1)}+E_{44}^{(1)}\}\ ,$
$\mathcal{B}_{2}$ $=\{E_{41}^{(0)},E_{42}^{(0)},E_{43}^{(1)},E_{43}^{(0)}\}\ .$
For $\in\mathcal{B}_{1}, aa$ and $J_{a}$ are generators of $U(\lambda,\mu)$ and $W^{k}(\lambda,\mu)$, respectively, of conformal weight $1$. The generators of $U(\lambda,\mu)$ of conformal weight $2$ are
@@
If $\lambda=(1,1,2,3)$ and $\mu=(1,1,2)$, then the last terms of (5.7) and (5.8) are equal to 0. In this case, the generators of $W^{k}(\lambda,\mu)$ of conformal weight $2$ are